# HyperSBINN: A Hypernetwork-Enhanced Systems Biology-Informed Neural Network for Efficient Drug Cardiosafety Assessment

### v1

@@ Abstract
Keywords: System-Biology informed neural network, Hypernetwork and Cardiac action potential.
@@ 1 Introduction
To solve this problem, this study introduces the use of HyperPINN in conjunction with a highly validated mathematical CAP model presented by Mohr et al. (2022). HyperPINN, developed by de Avila Belbute-Peres et al. (2021), adopts a meta-learning approach, utilizing a separate network (hypernetwork) to understand how parameters influence the equation. This allows the primary network responsible for solving the equation to remain simple and efficient.
In our specific case, the parametrization variables of the ODE systems pertain to the compound’s half-maximal inhibitory concentration (IC50) for a selection of four ionic channels: hERG potassium channel, hKv4.3 potassium channel, hCav1.2 L-type calcium channel, and hNav1.5 sodium channel. Each set of these four parameters characterizes the inhibitory potency of an individual compound on these selected channels. Given our objective of testing multiple compounds simultaneously, we encounter challenges in parametrized physics-informed neural networks, such as solving complex differential equations with a high number of parameters, handling high-dimensional parameters space, and ensuring accurate predictions across various parameter combinations .
This paper presents results utilizing HyperPINN, referred to as HyperSBINN, integrating hypernetworks and system biology-informed neural networks (SBINNs). Our approach highlights the significance and effectiveness of employing hypernetworks to train the parameter space of IC50 values. We compare our findings with other machine learning techniques, to demonstrate the efficacy of HyperSBINN.
@@ 2 Cardiac action potential model
Our study focuses on the modelling of the cardiac action potential (CAP) under different drug configurations. We focus on the model proposed by Mohr et al. (2022), which we briefly recall here.
In our context, the effect of a drug on CAP is supposed to depend only on its concentration (denoted $\mathsf{D}$) and on the inhibitory effect of the drug on $J$ ionic channels. The inhibitory effect of the $j$-th ($1\leqslant j\leqslant J$) ionic channel is denoted by $\mathsf{IC50}_{j}$. We denote $\mathsf{X}=\{\mathsf{D},\mathsf{IC50}_{1},\dots,\mathsf{IC50}_{J}\}$ the controlled characteristics of the drug. The CAP $V(t,\mathsf{X})$ at time $t$ under the configuration $\mathsf{X}$ is related to $J=8$ ionic currents (listed in the Appendix A), denoted by $I_{j}(t,\mathsf{X})$ ($1\leqslant j\leqslant 8$) through the ordinary differential equation (ODE)  [^1] This is the classical equation for electrical properties of a single cardiac cell, where we set the cell capacitance to 1, as in Mohr et al. (2022).
@@
where $I_{\text{stim}}(t)$ is a known externally applied stimulus current (see Appendix). The generic formulation of an ionic current $I_{j}(t,\mathsf{X})$ from a channel $j$ is based on Ohm’s law and can be stated as
$I_{j}(t,\mathsf{X})=\mathsf{G}_{j}\cdot k_{j}(\mathsf{D},\mathsf{IC50}_{j})\cdot g_{j}\left(t,V(t,\mathsf{X})\right)\cdot(V(t,\mathsf{X})-\mathsf{E}_{j})\,,$ (2)
@@
  • $k_{j}(\mathsf{D},\mathsf{IC50}_{j})$ is a function describing the inhibition factor for channel $j$. This function is either always 1 (no inhibition), or takes the following form for the four ionic channels listed in the Appendix:
    $k_{j}(\mathsf{D},\mathsf{IC50}_{j})=\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{j}}\right)^{-1}\,.$
  • $g_{j}(t,V(t,\mathsf{X}))$ is the fraction of conducting channels, which depends on voltage and a set of kinetic parameters. It depends on $V(t,\mathsf{X})$ through ODEs (resulting in 13 ODEs, all listed in the Appendix A).
The complexity of the proposed models lies in the expressions of $V(t,\mathsf{X})$ and the different $g_{j}(t,V(t,\mathsf{X}))$, which requires solving a (nonlinear) system of 14 ODEs, whose solution has a nonlinear dependence on $\mathsf{X}$, to obtain a CAP curve for each drug configuration. We denote by $u(t,\mathsf{X})$ the function taking values in $\mathbb{R}^{14}$ which is the solution of our system of ODE. The next section depicts the framework to obtain a fast approximation of this function for every $t$ and $\mathsf{X}$ without using an ODE solver for each configuration $\mathsf{X}$.
@@ 3.1 Overview of Physics-Informed Neural Networks
In this first part, we focus on the approximation of the solution for a given $\mathsf{X}$. Thus this variable is omitted in the notation. Consider the problem of approximating a function $u(t)\in\mathbb{R}^{d}$, for $t\in[0,T]$ (where $T$ is a fixed time horizon), such that $u(t)$ is the solution of a nonlinear system of differential equations, expressed as:
$\begin{array}[]{rl}u(0)&=\mathsf{u_{0}}\\ \frac{{\rm{d}}u}{{\rm{d}}t}(t)&=f(u(t))\,,\end{array}$ (3)
where $f(\cdot)$ is a known non linear function and $\mathsf{u_{0}}$ is the known initial condition. We suppose moreover that we have a set of measurements $y_{1:n}:=y_{1},\dots,y_{n}$ from the target function [^2] Still, in this section, observations would come from a single drug configuration $\mathsf{X}$. at times $t_{1:n}:=t_{1},\dots,t_{n}$
Physics-informed neural networks (PINNs, Raissi et al., 2019), approximates the solution $u$ by a function $v_{\text{}}(t,\theta)$, which is a neural network parameterized by a set of weights and biases $\theta$. Neural networks provide a flexible set of arbitrarily complex functions that performs well in approximating non linear functions and whose derivatives (with respect to $t$ or $\theta$) can be easily computed by automatic differentiation in standard machine learning softwares. For a given neural network architecture [^3] Defined by the number of layers, the number of neurons per layers, the connections between layers, and the activation functions., the approximation problem boils down to learn $\theta$ that minimize some loss function $\mathcal{L}^{\text{}}(\theta)$. In our context, a ”good” approximation is a function $v_{\text{}}(t,\theta)$ such that:
@@
Note that the choice of the weights $w^{\text{Data}},w^{\text{IC}}$ and $w^{\text{ODE}}$ will affect the training. This choice is model-driven and is discussed for our case study in Section 5.2. Note that we wrote Equation (3) with a common weight for all ODEs, but it is straightforward to have different weights for the different ODEs (which turned out to be necessary in our case-study).
@@ 3.2 HyperPINN
where $\mathcal{L}^{\text{PINN}}$ is defined in Equation (4). Figure 1 depicts graphically the framework of hyperPINNs.
@@
Figure 1: The two-stage neural network architecture of HyperPINNs. The hypernetwork takes a set of configuration $\mathsf{X}$ as input and outputs the weights and biases for the main PINNs network. The PINNs network then takes a time coordinate as input and makes predictions based on the problem defined by the hypernetwork output.
@@ 4 Experimental settings
In our approach, called HyperSBINN, we combine systems biology-informed neural networks, defined in Yazdani et al. (2020), with hypernetworks to predict CAP at various concentrations. Our input $\mathsf{X}$ is a vector of size $5$, stacking the IC50 values of four cardiac ion channel targets of interest (hERG, hKv3.4, hCav1.2, and hNav1.5), and the drug concentration. The model then outputs 14 curves, and in particular the curve of $V(t)$, which is our main target.
We assess the performance of our approach by first comparing the curves outputted by the HyperSBINN to the curves outputted by the solver and second, by computing a quantity of interest (biomarker) from the voltage curve, namely APD90. Since this quantity of interest is scalar, we can compare our HyperSBINN approach with standard machine learning techniques.
@@ 4.1 Assessing performance of the hyperSBINN model
We asses the performance of the hyperSBINN model to retrieve the solution of the ODE system for different set of drug configurations when the train is performed using our different losses. As observations, we provide outputs of a SBINN solver at random times (uniformly in [0, 500ms]) for random drug configurations. we uniformly randomly generated a total of 500 different drug configurations. These configurations include uniformly simulated IC50 values for four ionic channels, ranging from $0.1\,\mu M$ to $100\,\mu M$ , and drug concentrations $\mathsf{D}$ varying between $0\,\mu M$ and $4\,\mu M$.
@@ 4.2 Learning specific biomarker of the CAP
As mentioned previously, the curve $V(t)$ is the target output. From this curve, we can extract significant biomarker responses such as the APD50, the peak value, and others. An important biomarker is the APD90 response, which is defined as the time it takes for the membrane potential to repolarize to $90\%$ of its peak value during an action potential. Mathematically, it can be expressed as:
[…]
where $t_{90}$ is the time at which the membrane potential $V(t)$ repolarizes to $90\%$ of its peak value, and $t_{peak}$ is the time at which $V(t)$ reaches its peak value. These values are obtained directly from the $V(t)$ curves. To evaluate the effectiveness of the HyperSBINN model, we will compare its predictions of APD90 values across a set of the same drug configurations as in the previous part. The HyperSBINN model was used to predict the voltage curves $V(t)$, from which the APD90 values were subsequently extracted. The absolute difference with the numerical solver (considered as the ground truth) is then computed for each configuration in the test set. Additionally, we will compare the APD90 predicted by the HyperSBINN with those obtained from a Random Forest model, a standard machine learning technique which has provided the best performances in this case. While the methodologies differ, the comparison aims to evaluate performance of predicting the APD90. The Random Forest model will be trained using the same configuration set as the hyperSBINN (see Section 2), and will predict APD90 values directly, bypassing the voltage curve $V(t)$ and the CAP model.
@@ 5 Implementation details
In this section, we provide details of our implementation HyperSBINN. Python (3.10) was used for all numeric computations. HyperSBINN model was designed, trained and deployed using the Python library Jinns  [^4] https://github.com/mia-jinns/jinns. Note that the training of the HyperSBINN is time demanding, its evaluation (once trained) is fast.
@@ 5.1 HyperSBINN architecture and training
Regarding the activation function, we use the hyperbolic tangent activation function for all layers except the last one. The output layer matches the possible values taken by the solutions of the ODEs, which is [0, 1] for the 13 gate functions and ];+[]$\infty$;$\infty$[] - ∞ ; + ∞ [ for the voltage function. The match is made by using sigmoid and linear activation functions respectively. For the hypernetwork, we consistently use the hyperbolic tangent activation function.
@@ 5.2 Losses weights
The weight $w^{\text{IC}}$ for the initial condition loss term was set to 1 to prioritize satisfying the initial points during training, ensuring the model adheres to the starting state of the system. Additionally, the weights for the observational data loss $w^{\text{Data}}$ were also set to 1. This configuration allows the model to learn from the physical dynamics and governing ODEs while still incorporating observations, but without giving them undue influence over the training process.
5.3 Collocation points
To avoid overfitting to specific collocation points, the ODE loss is computed with batches of 500 times (in [0, 500]) which where uniformly sampled at each iteration. This sampling strategy certainly avoids overfitting, but could be improved by sampling in times zones where the model performs poorly.
For readers who want to reproduce our study, it is worth noting that a preprocessing step is necessary to ensure efficient neural network training. We address this by employing a linear scaling mechanism for the time variable $t$. This involves normalization by dividing $t$ by the maximum time domain value, $T_{max}$. This transformation $t=t/T_{max}$ compresses the $t$ scale to a more manageable range, typically around O(1), for optimal NN performance.
5.4 Observational Data points
As we stated in Section 4, we will add some observational data ($y_{1:n}$) to the network. We derive these observations by extracting configurations $\mathsf{X}$ from the hypernetwork, and then we used a numerical solver to generate the solution of the CAP model given these configurations. For the numerical solver, we use the same initial conditions provided to the HyperSBINN model and described above.
Generated data are provided to the network as an input and used to minimise the data discrepancy loss function, and introduced in the loss function $l_{Data}$ in Equation (4). From the extensive dataset, we randomly select 0.6 % data points of the total generated data, injecting them into the hyperSBINN along with their corresponding time and configuration $\mathsf{X}$. This augmentation strategy enhances the robustness and generalization capabilities of our network architecture.

### v2

@@ Abstract
Keywords: System-Biology informed neural network, hypernetwork and Cardiac action potential
@@ 1 Introduction
This study introduces a third approach, named hyperSBINN (for hyper System biology informed neural network). This method mimicks hyperPINN, developed by de Avila Belbute-Peres et al. (2021), which adopts a meta-learning approach, utilizing a separate network (hypernetwork) to understand how parameters influence the equation. This allows the primary network responsible for solving the equation to remain simple and efficient.
When PDEs or ODEs are parameterized by a function that specifies initial or boundary conditions (the input function), operator learning methods—such as deep operator networks or Fourier neural operators, which learn the mapping from the input function to the solution of the PDE or ODE—have been applied in systems biology with success (Li et al., 2024; Centofanti et al., 2024; Pellegrini et al., 2025). In contrast, our system differs in that we use scalar parameters to parameterize the ODEs. Given that the system under study assumes consistent initial conditions and involves a scalar parametrization, we focus on the hyperSBINN approach, even though an operator learning method could also accommodate scalar parameters (see Huang et al. (2024), for instance.
Our method is illustrated on a highly validated mathematical CAP model presented by Mohr et al. (2022). The parametrization variables of the ODE systems pertain to the compound’s half-maximal inhibitory concentration (IC50) for a selection of four ionic channels: hERG potassium channel, hKv4.3 potassium channel, hCav1.2 L-type calcium channel, and hNav1.5 sodium channel. Each set of these four parameters characterizes the inhibitory potency of an individual compound on these selected channels. Given our objective of testing multiple compounds simultaneously, we encounter challenges in parametrized physics-informed neural networks, such as solving complex differential equations with a high number of parameters, handling high-dimensional parameters space, and ensuring accurate predictions across various parameter combinations .
Our approach highlights the significance and effectiveness of employing hypernetworks to train the parameter space of IC50 values. We compare our findings with other machine learning techniques, to demonstrate the efficacy of hyperSBINN.
@@ 2 Cardiac action potential model
Our study focuses on the modelling of the cardiac action potential (CAP) under different drug configurations. We focus on the model proposed by Mohr et al. (2022), which we briefly recall here, and fully detail in Appendix A.
In our context, the effect of a drug on CAP is supposed to depend only on its concentration (denoted $\mathsf{D}$) and on the inhibitory effect of the drug on four ionic channels. The inhibitory effect of the $k$-th ($1\leqslant k\leqslant 4$) ionic channel is denoted by $\mathsf{IC50}_{k}$. We denote $\mathsf{X}=\{\mathsf{D},\mathsf{IC50}_{1},\dots,\mathsf{IC50}_{4}\}$ the five controlled characteristics of the drug. The CAP $V(t,\mathsf{X})$ at time $t$ under the configuration $\mathsf{X}$ is related to $J=8$ ionic currents (listed in the Appendix A), denoted by $I_{j}(t,\mathsf{X})$ ($1\leqslant j\leqslant 8$) through the ordinary differential equation (ODE)  [^1] This is the classical equation for electrical properties of a single cardiac cell, where we set the cell capacitance to 1, as in Mohr et al. (2022).
@@
where $I_{\text{stim}}(t)$ is a known externally applied stimulus current. The generic formulation of an ionic current $I_{j}(t,\mathsf{X})$ from a channel $j$ is based on Ohm’s law and can be stated as
$I_{j}(t,\mathsf{X})=\mathsf{G}_{j}\cdot k_{j}(\mathsf{D},\mathsf{IC50}_{k})\cdot g_{j}\left(t,V(t,\mathsf{X})\right)\cdot(V(t,\mathsf{X})-\mathsf{E}_{j})\,$, (2)
@@
  • $k_{j}(\mathsf{D},\mathsf{IC50}_{k})$ is a function describing the inhibition factor for channel $j$. This function is either always 1 (no inhibition), or takes the following form for the four ionic channels listed in the Appendix:
    $k_{j}(\mathsf{D},\mathsf{IC50}_{k})=\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{k}}\right)^{-1}\,.$
  • $g_{j}(t,V(t,\mathsf{X}))$ is the fraction of conducting channels, which depends on voltage and a set of kinetic parameters. All functions $g_{j}$ are defined thanks to 13 solutions of ordinary differential equations depending on $V(t,\mathsf{X})$. The 13 solutions of these ODEs lies in the interval [0, 1] and are referred to as gate functions.
The complexity of the proposed models lies in the expressions of $V(t,\mathsf{X})$ and the different $g_{j}(t,V(t,\mathsf{X}))$, which requires solving a (nonlinear) system of 14 ODEs (given by Equations (1) and (5)-(17), all listed in the Appendix A), whose solution has a nonlinear dependence on $\mathsf{X}$. We denote by $u(t,\mathsf{X})$ the function taking values in $\mathbb{R}^{14}$ which is the solution of our system of ODEs. The next section depicts the framework to obtain a fast approximation of this function for every $t$ and $\mathsf{X}$ without using an ODE solver for each configuration $\mathsf{X}$.
@@ 3.1 Overview of Physics-Informed Neural Networks
In this first part, we focus on the approximation of the solution for a given $\mathsf{X}$ (thus this variable is omitted in the notations). Consider the problem of approximating a function $u(t)\in\mathbb{R}^{d}$, for $t\in[0,T]$ (where $T$ is a fixed time horizon), such that $u(t)$ is the solution of a nonlinear system of differential equations, expressed as:
[…]
where $f(\cdot)$ is a known non linear function and $\mathsf{u_{0}}$ is the known initial condition. We suppose moreover that we have a set of measurements $y_{1:n}:=y_{1},\dots,y_{n}$ from the target function [^2] Still, in this section, observations would come from a single drug configuration $\mathsf{X}$. at times $t_{1:n}:=t_{1},\dots,t_{n}$ .
Physics-informed neural networks (PINNs, Raissi et al., 2019), approximates the solution $u$ by a function $v_{\text{}}(t,\theta)$, which is a neural network parameterized by a set of weights and biases $\theta$. Neural networks provide a flexible set of arbitrarily complex functions that performs well in approximating non linear functions and whose derivatives (with respect to $t$ or $\theta$) can be easily computed by automatic differentiation in standard machine learning softwares. For a given neural network architecture [^3] Defined by the number of layers, the number of neurons per layers, the connections between layers, and the activation functions., the approximation problem boils down to learn $\theta$ that minimized some loss function $\mathcal{L}^{\text{}}(\theta)$. In our context, a ”good” approximation is a function $v_{\text{}}(t,\theta)$ such that:
@@
Note that the choice of the weights $w^{\text{Data}},w^{\text{IC}}$ and $w^{\text{ODE}}$ will affect the training. This choice is model-driven and is discussed for our case-study in Section 5.4. Moreover, for sake of simplicity, we wrote the ODE loss with a common weight for all of its 14 components (given by our CAP model), but it is straightforward to have different weights for the different ODEs (which turned out to be necessary in our case-study).
@@ 3.2 HyperPINN
where $\mathcal{L}^{\text{PINN}}$ is defined in Equation (4). As a hypernetwork, a hyperPINN is expected to have fewer parameters (weights) to learn than the corresponding dense neural network. This should then lead to a faster training (Chauhan et al., 2024). Because the ODE system represents a biological process, we refer to the hyperPINN as a hyperSBINN in the following. Once trained, the hyperSBINN acts as a quasi-instantaneous solver of the ODE system for any input configuration $\mathsf{X}$, effectively serving as a meta-model by replacing the numerical solver with a faster alternative. Figure 1 depicts graphically the framework of hyperSBINNs, highlighting that the architecture relies on two neural networks: the hypernetwork that provides weights and biases for the SBINN. A parallel can be drawn with Przedborski et al. (2021), who also pursue meta-modeling; however, they employ a single neural network architecture and omit the inclusion of an ODE loss during training. Our meta-modeling approach differs from that of Yazdani et al. (2020) and Daneker et al. (2023), whose objective is to estimate ODE parameters while simultaneously training the SBINN. In their case, the SBINNs provide solutions only in the neighborhood of the estimated parameters, rather than for any configuration.
@@
Figure 1: The two-stage neural network architecture of hyperSBINN. The hypernetwork takes a set of configuration $\mathsf{X}$ as input and outputs the weights and biases for the main PINNs network. The PINNs network then takes a time coordinate as input and makes predictions based on the problem defined by the hypernetwork output.
@@ 4 Experimental settings
We aim at training an hyperSBINN to predict CAP at various concentrations. Our input $\mathsf{X}$ is a vector of size $5$, stacking the IC50 values of four cardiac ion channel targets of interest and the drug concentration. The model then outputs 14 curves, and in particular the curve of $V(t)$, which is our main target.
We assess the performance of our approach by first comparing the curves outputted by the hyperSBINN to the curves outputted by the solver and second, by computing a quantity of interest (biomarker) from the voltage curve, namely APD90.
@@ 4.1 Assessing performance of the hyperSBINN model
To assess the performance of the hyperSBINN model to retrieve the solution of our ODE system, we compute a normalized root mean square error (NRMSE) over a testing set of curves obtained through an ODE solver. Our testing set consists of 500 solutions for 500 different drug configurations that were not seen during the training. These configurations include uniformly simulated IC50 values for four ionic channels, ranging from $0.1\,\mu M$ to $100\,\mu M$ , and drug concentrations $\mathsf{D}$ varying between $0\,\mu M$ and $4\,\mu M$. For each test configuration $\mathsf{X}_{i},(1\leq i\leq 500)$, the solver outputs a function $u^{*}(t,\mathsf{X})=\left(u^{*}_{j}(t,\mathsf{X})\right)_{1\leq j\leq 14}$ at testing times $t$ evenly spaced between $t_{0}=0$ and $t_{5000}=500$. We denote $\hat{u}_{j}(t_{k},\mathsf{X}_{i})$ the output of the hyperSBINN model for channel $j$, drug configuration $\mathsf{X}_{i}$ and time $t_{k}$. The NRMSE is then computed for each test configuration and each channel as follows:
$\mathsf{NRMSE}(i,j)=\frac{1}{r_{j}}\times\sqrt{\frac{1}{5001}\sum_{k=0}^{5000}\left(\hat{u}_{j}(t_{k},\mathsf{X}_{i})-u_{j}^{*}(t_{k},\mathsf{X}_{i})\right)^{2}}\,$,
where $r_{j}$ is a normalizing term to allow comparison between channels. Formally, $r_{j}$ is the spread of possible values for the $j$-th channel (1 for all channels but the voltage, whose spread is between -80 and 80). In purpose of comparison, we also compute this metric with three alternatives:
  • The same hyperSBINN architecture trained without the system biology losses. The training only relies on the training observations given by the solver.
  • The same hyperSBINN architecture trained without the observation loss. The training only relies on the system biology information, i.e., on the ODE and initial condition losses.
  • An SBINN without the hypernetwork structure similar to Przedborski et al. (2021). We train a single neural network whose inputs are given by $(\mathsf{X},t)$ and all the weights are to be learned. In the following, this method is referred to as metaSBINN.
@@ 4.2 Learning specific biomarker of the CAP
As mentioned previously, the curve $V(t)$ is the main target output. From this curve, we can extract significant biomarker responses such as the APD50, the peak value, and others. An important biomarker is the APD90 response, which is defined as the time it takes for the membrane potential to repolarize to $90\%$ of its peak value during an action potential. Mathematically, it can be expressed as:
$\text{APD90}=t_{90}-t_{\text{peak}},$
where $t_{90}$ is the time at which the membrane potential $V(t)$ repolarizes to $90\%$ of its peak value, and $t_{\text{peak}}$ is the time at which $V(t)$ reaches its peak value. These values are obtained directly from the $V(t)$ curves. To evaluate the effectiveness of the hyperSBINN model to predict this biomarker, we compare its predictions of APD90 values across a set of the same drug configurations as in the previous part. The hyperSBINN model was used to predict the voltage curves $V(t)$, from which the APD90 values were subsequently extracted. The absolute difference with the numerical solver (considered as the ground truth) is then computed for each configuration in the test set. Additionally, we will compare the APD90 predicted by the hyperSBINN with those obtained from a Random Forest model, a standard machine learning technique which has provided the best performances in this case. While the methodologies differ, the comparison aims to evaluate performance of predicting the APD90. The Random Forest model will be trained using the same training set as the hyperSBINN (see Section 2), and will predict APD90 values directly, bypassing the voltage curve $V(t)$ and the CAP model.
@@ 5 Implementation details
In this section, we provide details of our implementation hyperSBINN. Python 3.11 was used for all numeric computations. HyperSBINN model was designed, trained and deployed using the Python library jinns v1.6.1 (Gangloff and Jouvin, 2024) which uses JAX as a backend. Note that the training of the hyperSBINN is time demanding, its evaluation (once trained) is fast.
@@ 5.1 HyperSBINN architecture and training
Regarding the activation function, we use the hyperbolic tangent activation function for all layers except the last one. The output layer matches the possible values taken by the solutions of the ODEs, which is [0, 1] for the 13 gate functions and $\infty$$\infty$[ for the voltage function. The match is made by using sigmoid and linear activation functions respectively. For the hypernetwork, we consistently use the hyperbolic tangent activation function.
@@
5.2 Collocation points
The ODE loss is computed on a set of 200, 000 times sampled uniformly between 0 and 500, and 200, 000 drug configurations ($\mathsf{X}$). For the latter, the sampling is uniform in dimension 5, with the intervals being $[0.1,100]$ $\mu M$ for the IC50 values, and $[0;4]$ $\mu M$ for the drug concentration. This results in $4\times 10^{8}$ possible collocation points $(t,\mathsf{X})$. Each iteration of the stochastic gradient descent is performed on a mini batch of 500 collocation points.
For readers who want to reproduce our study, it is worth noting that a preprocessing step is necessary to ensure efficient neural network training. We address this by employing a linear scaling mechanism for the time variable $t$. This involves normalization by dividing $t$ by the maximum value of the time domain, $T_{max}$. This transformation $t=t/T_{max}$ compresses the $t$ scale to a more manageable range, typically around O(1), for optimal NN performance.
5.3 Synthetic data
The synthetic data used to compute the observation loss are outputs of a Livermore solver for ordinary differential equations (Radhakrishnan and Hindmarsh, 1993). The solution of the ODE is obtained for 500 different IC50 and drug configurations, sampled uniformly in the same space as the test set or the collocation points (Sections 4.1 and 5.2). The solver outputs the solution with time step 0.1 between $t=0$ and $t=100$, and with time step 0.5 between $t=100$ and $t=500$ (the beginning is more finely sampled since the important features of the signal tend to occur during this period). To avoid overfitting to specific time points, the solver’s solution is then evaluated at 100 times chosen uniformly at random.
@@ 5.4 Losses weights
The weight $w^{\text{IC}}$ for the initial condition loss term was set to 10 to prioritize satisfying the initial points during training, ensuring the model adheres to the starting state of the system. Additionally, the weights for the observational data loss $w^{\text{Data}}$ were set to 1. This configuration allows the model to learn from the physical dynamics and governing ODEs while still incorporating observations, but without giving them undue influence over the training process. As described in Section 4.1, we consider two alternatives of the hyperSBINN where, on one hand, the weights $w^{\text{IC}}=w^{\text{ODE}}=0$, leading to a model learning without biological information (and thus only based on synthetic data), and, on the other hand, $w^{\text{Data}}=0$, leading to a model learning without synthetic data (and thus only biology informed).